# A Survey on Sparse Autoencoders: Interpreting the Internal Mechanisms of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionized natural language processing, yet their internal mechanisms remain largely opaque. Recently, mechanistic interpretability has attracted significant attention from the research community as a means to understand the inner workings of LLMs. Among various mechanistic interpretability approaches, Sparse Autoencoders (SAEs) have emerged as a particularly promising method due to their ability to disentangle the complex, superimposed features within LLMs into more interpretable components. This paper presents a comprehensive examination of SAEs as a promising approach to interpreting and understanding LLMs. We provide a systematic overview of SAE principles, architectures, and applications specifically tailored for LLM analysis, covering theoretical foundations, implementation strategies, and recent developments in sparsity mechanisms. We also explore how SAEs can be leveraged to explain the internal workings of LLMs, steer model behaviors in desired directions, and develop more transparent training methodologies for future models. Despite the challenges that remain around SAE implementation and scaling, they continue to provide valuable tools for understanding the internal mechanisms of large language models.
@@ A Survey on Sparse Autoencoders:
[…]
Figure 1: (a) This figure illustrates the fundamental framework of a Sparse Autoencoder (SAE). SAE is trained to take a model representation $\mathbf{z}$ as input and project it to an overcomplete sparse activation $h(\mathbf{z})$ by learning to reconstruct the original input $\hat{\mathbf{z}}$. The SAE typically comprises an encoder, a decoder, and a loss function for training. (b) The development of the SAE progresses through multiple stages, with each stage drawing inspiration from and building upon the previous one.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI et al. (2024), Claude-3.5 Anthropic (2024), DeepSeek-R1 DeepSeek-AI et al. (2025), and Grok-3 xAI (2025), have emerged as powerful tools in natural language processing, demonstrating remarkable capabilities in tasks ranging from text generation to complex reasoning. However, their increasing size and complexity have created significant challenges in understanding their internal representations and decision-making processes. This “black box” nature of LLMs has sparked growing interest in mechanistic interpretability Bereska and Gavves (2024); Zhao et al. (2024a); Rai et al. (2024); Zhao et al. (2024b), a field that aims to break down LLMs into understandable components and systematically analyze how these components interact to produce emergent behaviors. Understanding these mechanisms is crucial not only for scientific progress but also for addressing concerns related to safety, reliability, and alignment of increasingly powerful LLM systems.
@@
Among the various approaches to interpreting LLMs, Sparse Autoencoders (SAEs) (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2025; Rajamanoharan et al., 2024b) have emerged as a particularly promising direction for addressing a fundamental challenge in LLM interpretability: polysemanticity. Many neurons in LLMs are polysemantic, responding to seemingly unrelated concepts or features simultaneously. This is a phenomenon likely resulting from superposition Elhage et al. (2022), where LLMs represent more independent features than they have neurons by encoding each feature as a linear combination of neurons. SAEs address this issue by learning an overcomplete, sparse representation of neural activations, effectively disentangling these superimposed features into more interpretable units. By training a sparse autoencoder to reconstruct the activations of a target network layer while enforcing sparsity constraints, SAEs can extract a larger set of monosemantic features that offer clearer insights into what information the LLM is processing. This approach has shown considerable promise in transforming the often-inscrutable activations of LLMs into more human-understandable representations, potentially creating a more effective vocabulary for mechanistic analysis of these complex systems.
@@
In this paper, we provide a comprehensive overview of SAE for LLM interpretability, beginning with their foundational motivations and development history. We then explore the technical framework of SAEs, including their basic architecture, various design improvements, and effective training strategies. The survey examines different approaches to analyzing and explaining SAE features, categorized broadly into input-based and output-based explanation methods. We discuss evaluation methodologies for assessing SAE performance, covering both structural metrics that analyze the properties of the learned features and functional metrics that evaluate their practical utility. The paper further delves into real-world applications of SAEs in understanding and manipulating LLMs, including model behavior analysis, intentional steering of model outputs, and insights for improved model training. Lastly, we highlight current research challenges and conclude with perspectives on promising future directions.
2 Why Sparse Autoencoders?
Explaining Individual Neurons.
The development of interpretability techniques for LLMs has progressed in stages rather than as a single step. From 2022 to 2023, researchers at OpenAI and Anthropic focused on understanding LLMs by examining individual neurons. OpenAI, for instance, leveraged GPT-4 to generate natural language explanations for neurons in models like GPT-2, attempting to map specific neuron activations to concrete linguistic or conceptual features Bills et al. (2023). Similarly, Anthropic built small toy models trained on synthetic data to observe how features are stored in neurons. These early investigations showed that analyzing single neurons can provide initial insights Elhage et al. (2022). In addition, it is worth noting that the study of explaining individual neurons by labeling interpretable features to them has been extensively explored in non-MI studies (Radford et al., 2017; Donnelly and Roegiest, 2019; Nguyen et al., 2019; Szegedy et al., 2013) prior to the introduction of the MI.
SAEs for Small-size Language Models.
SAEs for Large Language Models.
@@ 2 Why Sparse Autoencoders?
Improve Architecture (§3.2.1) Gated SAE Jump ReLU Rajamanoharan et al. (2024a)
@@
Improve Training Strategy (§3.2.2) Layer Group SAE Jump ReLU Ghilardi et al. (2024)
@@ 3.1 Basic SAE Framework
SAE is a neural network that learns an overcomplete dictionary for representation reconstruction. As shown in Figure 1a, the input of SAE is the representation of a token extracted from LLMs, which is mapped onto a sparse vector of dictionary activations.
@@
Encoding Step.   Given an input representation $\mathbf{z}\in\mathbb{R}^{d}$, the encoder applies a linear transformation using a weight matrix $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{d\times m}$ and a bias term $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{m}$, followed by a ReLU activation function to enforce sparsity. The encoding operation is defined as:
$h(\mathbf{z})=\text{ReLU}\left(\mathbf{z}\cdot\mathbf{W}_{\text{enc}}+\mathbf{b}_{\text{enc}}\right),$ (1)
where $h(\mathbf{z})\in\mathbb{R}^{m}$ represents the sparse activation vector, which helps disentangle superposition features. The ReLU activation function $\text{ReLU}\left(x\right)=\max\left(0,x\right)$ ensures that only non-negative values pass through, which encourages sparsity by setting negative values to zero.
@@
where $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{m\times d}$ is the decoder weight matrix. $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$ is the decoder bias term. $\hat{z}\in\mathbb{R}^{d}$ is the reconstructed output, which aims to approximate the original input $\mathbf{z}$.
@@
Loss Function. The activation vector $h(\mathbf{z})$ is encouraged to be sparse, meaning that most of its values should be zero. While the ReLU activation function after the encoder enforces basic sparsity by setting negative values to zero, it does not necessarily eliminate small positive values, which can still contribute to a dense representation. Therefore, additional sparsity enforcement is required. This is achieved using a sparsity regularization term in the loss function, which further promotes a minimal number of active features. Beyond enforcing sparsity, the SAE must also ensure that the learned sparse activation retains sufficient information to accurately reconstruct the original input $\mathbf{z}$. The loss function for training the SAE consists of two key components: reconstruction loss and sparsity regularization:
@@ 3.2 Different SAE Variants
As SAEs continue to emerge as a powerful tool for interpreting the internal representations of LLMs, researchers have increasingly focused on refining and extending their capabilities. Various SAE variants have been proposed to address the limitations of traditional SAEs, each introducing improvements from different perspectives. In this section, we categorize these advancements into two main groups: Improve Architectural, which modify the structure and design of the traditional SAE, and Improve Training Strategy, which retain the original architecture but introduce novel methods to enhance training efficiency, feature selection, and sparsity enforcement.
3.2.1 Improve Architecture

### v2

@@ Abstract
Large Language Models (LLMs) have transformed natural language processing, yet their internal mechanisms remain largely opaque. Recently, mechanistic interpretability has attracted significant attention from the research community as a means to understand the inner workings of LLMs. Among various mechanistic interpretability approaches, Sparse Autoencoders (SAEs) have emerged as a promising method due to their ability to disentangle the complex, superimposed features within LLMs into more interpretable components. This paper presents a comprehensive survey of SAEs for interpreting and understanding the internal workings of LLMs. Our major contributions include: (1) exploring the technical framework of SAEs, covering basic architecture, design improvements, and effective training strategies; (2) examining different approaches to explaining SAE features, categorized into input-based and output-based explanation methods; (3) discussing evaluation methods for assessing SAE performance, covering both structural and functional metrics; and (4) investigating real-world applications of SAEs in understanding and manipulating LLM behaviors.
@@ A Survey on Sparse Autoencoders:
Figure 1: (a) This figure illustrates the fundamental framework of a Sparse Autoencoder (SAE). SAE is trained to take a model representation $\mathbf{z}$ as input and project it to an overcomplete sparse activation $h(\mathbf{z})$ by learning to reconstruct the original input $\hat{\mathbf{z}}$. The SAE typically comprises an encoder, a decoder, and a loss function for training. (b) The development of the SAE progresses through multiple stages. Note that we only list some representative SAE models in this timeline rather than providing an exhaustive compilation.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI et al. (2024), Claude-3.5 Anthropic (2024), DeepSeek-R1 DeepSeek-AI et al. (2025), and Grok-3 xAI (2025), have emerged as powerful tools in natural language processing, demonstrating remarkable capabilities in tasks ranging from text generation to complex reasoning. However, their increasing size and complexity have created significant challenges in understanding their internal representations and decision-making processes. This “black box” nature of LLMs has sparked a growing interest in mechanistic interpretability Bereska and Gavves (2024a); Zhao et al. (2024a); Rai et al. (2024); Zhao et al. (2024b), a field that aims to break down LLMs into understandable components and systematically analyze how these components interact to understand their behaviors.
@@
Among the various approaches to interpreting LLMs, Sparse Autoencoders (SAEs)  (Cunningham et al., 2023; Bricken et al., 2023; Gao et al., 2025; Rajamanoharan et al., 2024b; Galichin et al., 2025) have emerged as a particularly promising direction for addressing a fundamental challenge in LLM interpretability: polysemanticity. Many neurons in LLMs are polysemantic, responding to seemingly unrelated concepts or features simultaneously. This is a phenomenon likely resulting from superposition Elhage et al. (2022), where LLMs represent more independent features than they have neurons by encoding each feature as a linear combination of neurons. SAEs address this issue by learning an overcomplete, sparse representation of neural activations, effectively disentangling these superimposed features into more interpretable units. By training a sparse autoencoder to reconstruct the activations of a target network layer while enforcing sparsity constraints, SAEs can extract a larger set of monosemantic features that offer clearer insights into what information the LLM is processing. This approach has shown promise in transforming the often-inscrutable activations of LLMs into more human-understandable representations, potentially creating a more effective vocabulary for mechanistic analysis of these complex systems.
1.1 Contribution and Uniqueness
@@
Our Contributions. In this paper, we provide a comprehensive overview of SAE for LLM interpretability, with major contributions listed as following: (1) We explore the technical framework of SAEs, including their basic architecture, various design improvements, and effective training strategies (Section 2). (2) We examine different approaches to analyzing and explaining SAE features, categorized broadly into input-based and output-based explanation methods (Section 3). (3) We discuss evaluation methodologies for assessing SAE performance, covering both structural metrics and functional metrics (Section 4). (5) We discuss real-world applications of SAEs in understanding and manipulating LLMs (Section 5). (6) Additionally, in the appendix, we also introduce key motivations for SAE, discuss connection of SAEs to the broader field of mechanistic interpretability, provide experimental evaluations, and highlight current research challenges and promising future directions.
Differences with Existing Surveys. Several existing surveys take a broad perspective on LLM interpretability. For instance, some surveys provide comprehensive overviews of general explainability methods for LLMs Ferrando et al. (2024); Zhao et al. (2024a), while others focus specifically on mechanistic interpretability as a whole Rai et al. (2024); Bereska and Gavves (2024b). In contrast, our work uniquely focuses exclusively on SAEs as a specific and promising approach within the mechanistic interpretability landscape. By narrowing our scope to SAEs, we are able to provide a much more comprehensive and detailed analysis of their principles, architectures, training methodologies, evaluation techniques, and practical applications.
@@
Improve Architecture (§C.1) Gated SAE Jump ReLU Rajamanoharan et al. (2024a)
@@
Improve Training Strategy (§C.2) Layer Group SAE Jump ReLU Ghilardi et al. (2024)
@@ 2.1 Basic SAE Framework
SAE is a neural network that learns an overcomplete dictionary for representation reconstruction. As shown in Figure 1a, the input of SAE is the representation of a token from LLMs, which is mapped onto a sparse vector of dictionary activations.
@@
Encoding Step.   Given an input representation $\mathbf{z}\in\mathbb{R}^{d}$, the encoder applies a linear transformation using a weight matrix $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{d\times m}$ and a bias term $\mathbf{b}_{\text{enc}}\in\mathbb{R}^{m}$, followed by an activation function $\sigma$ to enforce sparsity. The encoding operation is defined as:
$h(\mathbf{z})=\sigma\left(\mathbf{z}\cdot\mathbf{W}_{\text{enc}}+\mathbf{b}_{\text{enc}}\right),$ (1)
where $h(\mathbf{z})\in\mathbb{R}^{m}$ represents the sparse activation vector, which helps disentangle superposition features. The $\sigma$ activation function could take different formats (see Table 1). For example, $\sigma$ could be $\text{ReLU}\left(x\right)=\max\left(0,x\right)$, ensures that only non-negative values pass through, encouraging sparsity by setting negative values to zero.
@@
where $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{m\times d}$ is the decoder weight matrix. $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$ is the decoder bias term. $\hat{\mathbf{z}}\in\mathbb{R}^{d}$ is the reconstructed output, which aims to approximate the original input $\mathbf{z}$.
@@
Loss Function. The activation vector $h(\mathbf{z})$ is encouraged to be sparse, meaning that most of its values should be zero. Take the ReLU activation for example, while the activation function after the encoder enforces basic sparsity by setting negative values to zero, it does not necessarily eliminate small positive values, which can still contribute to a dense representation. Therefore, additional sparsity enforcement is required. This is achieved using a sparsity regularization term in the loss function, which further promotes a minimal number of active features. Beyond enforcing sparsity, the SAE must also ensure that the learned sparse activation retains sufficient information to accurately reconstruct the original input $\mathbf{z}$. The loss function for training the SAE consists of two key components: reconstruction loss and sparsity regularization:
@@ 2.2 Different SAE Variants
As SAEs continue to emerge as a powerful tool for interpreting the internal representations of LLMs, researchers have increasingly focused on refining and extending their capabilities. Various SAE variants have been proposed to address the limitations of traditional SAEs, each introducing improvements from different perspectives. In this section, we categorize these advancements into two main groups: Improve Architectural, which modify the structure and design of the traditional SAE, and Improve Training Strategy, which retain the original architecture but introduce novel methods to enhance training efficiency, feature selection, and sparsity enforcement. A taxonomy of representative SAE frameworks is presented in Table 1. Due to page limitations, examples for each group are provided in Appendix C.1 (Improved Architectural) and Appendix C.2 (Improved Training Strategy). We also discuss challenges encountered during SAE training in Appendix C.3.
3 Explainability Analysis of SAEs
3.1 Input-based Explanations
where $f_{<l}(x)$ indicates generating the hidden representation of input text $x$ at the $l$-th layer, and $l$ is the layer our SAE is trained for. This strategy is reasonable for interpreting weight vectors of SAEs because of the sparse nature of SAEs, which indicates that a learned feature vector should only be activated by a certain pattern/concept. Therefore, summarizing the most activated text spans can give us a clue to understanding the semantic meaning encoded by the learned feature vector.
[…]
3.2 Output-based Explanations
VocabProj.   Output-based explanations project the learned feature vectors to the output word embeddings of words to compute the activations. Mathematically, $f_{\text{out}}(w):\mathcal{V}\rightarrow\mathbb{R}^{d}$ denotes the output word embedding layer that returns the output embeddings of a word $w$, and we can collect the natural language explanations by:
This mapping process makes sense for decoder-only LLMs because the layers in such models share the same residual stream, enabling the representations in the intermediate layers to be linear correlated to the output word embeddings Nostalgebraist (2020). Recently, researchers Wu et al. (2025b); Gur-Arieh et al. (2025) find that output-based explanations show a stronger promise in interpreting and controlling LLM behaviors (i.e., generated texts) compared to the input-based ones.
[…]
Compared with VocabProj, that only considers $p(w|\mathbf{w}_{m})$, this mutual-information-driven objective highlights the need to normalize the raw activation with $p(\mathbf{w}_{m}|w)$. That is, if a word whose output embedding consistently activates various feature vectors, it loses specification for interpretation.
4 Evaluation Metrics and Methods
4.1 Structural Metrics
Structural metrics focus on assessing whether an SAE behaves as intended during training. SAEs are designed to optimize both reconstruction fidelity and sparsity, as these properties are explicitly enforced in the training loss. Therefore, natural evaluation metrics assess reconstruction accuracy and sparsity in the model’s learned representations. Reconstruction Fidelity.   The most fundamental way to evaluate reconstruction fidelity is through Mean Squared Error (MSE) and Cosine Similarity Ng et al. (2011), which directly compare the original activations with SAE-reconstructed activations. Additional metrics such as Fraction of Variance Unexplained (FVU) (also known as normalized loss) Gao et al. (2025) and Explained Variance Karvonen et al. (2024) measure how much variance in the original data is retained after SAE reconstructs. Beyond direct reconstruction comparisons, researchers also evaluate how SAEs affect the probability distribution of model outputs. Cross-Entropy Loss Shannon (1948) and KL Divergence Kullback and Leibler (1951) measure the shift in probability distributions when substituting original model activations with SAE-generated activations. If the SAE faithfully reconstructs activations, the probability distributions should remain similar. Similarly, Delta LM Loss Lieberum et al. (2024) quantifies the difference between the original language model loss and the loss incurred when replacing activations with those from the SAE. Another important aspect of reconstruction fidelity is magnitude preservation. The L2 Ratio Karvonen et al. (2024) compares the Euclidean norms of different activations to ensure that the SAE does not systematically alter activation magnitudes.
4.2 Functional Metrics
Moreover, we provide a comprehensive comparison of SAEs using both structural and functional metrics in Appendix D.
5 Applications in Large Language Models
5.1 Model Steering
Unlike supervised concept vectors, such as probing classifiers Belinkov (2022); Zhao et al. (2025); Jin et al. (2025), SAEs can simultaneously learn a large volume of concept vectors. The learned vectors can then be utilized to steer model behaviors in ways similar to supervised vectors (See Figure 3). A detailed comparison between SAEs and probing-based methods is provided in Appendix F.
As previously mentioned, SAE latents can be employed to produce steering vectors that control model outputs related to toxicity, sycophancy, refusal, and emotions. Other steering applications have also proven feasible. Recently, a study shows that SAE latents are able to capture instructions such as translations. These identified latents are effective in manipulating models to translate inputs according to specific instructions He et al. (2025). Another investigation focusing on semantic search demonstrates that SAEs can learn fine-grained semantic concepts at various levels of abstraction. These concept vectors can be used to steer models toward related semantic O’Neill et al. (2024). Alternatively, SAEs trained on biological datasets can provide biology-related features that enable the unlearning of biology-relevant knowledge with fewer side effects than existing techniques Farrell et al. (2024). Given that steering effects are generally challenging to control, SAE-TS further utilizes SAE latents to optimize steering vectors by measuring the changes in SAE feature activation caused by steering, thereby helping construct vectors that target relevant features while minimizing unwanted effects Chalnev et al. (2024). Moreover, explanations based on SAE latents risk prioritizing linguistic features over semantic meaning. Wu et al. (2025b) propose a novel approach that promotes diverse semantic explanations, which has been demonstrated to enhance model safety.
5.2 Model Behavior Analysis
A recent study utilizes SAEs to reveal the mechanism of hallucination, where entity recognition plays a pivot role in recalling facts. A direction distinguishing whether the model knows an entity is identified, which is usually used for hallucination refusal in chat models Ferrando et al. (2025). Some studies focus on interpreting how in-context learning (ICL) is performed within models. One study focuses on general ICL tasks, and task-related function vectors were successfully isolated Kharlapenko et al. (2024). Another study focuses on reinforcement learning (RL) tasks. Their experiment shows that an LLM’s internal representations are capable of capturing temporal difference errors and Q-values that are essential in RL computations Demircan et al. (2025). Besides, one study attempts to examine the working mechanism of instruction following. Their analysis on translation tasks shows that instructions are composed of multiple relevant concepts rather than individual ones He et al. (2025).
Moreover, SAEs have been employed to study behaviors related to toxicity, sycophancy, refusal, and emotions. One recent study shows that features captured by SAEs can be used to construct probes to classify cross-lingual toxicity Gallifant et al. (2025). By examining SAE latents that activate on anger-related tokens, researchers have identified steering vectors that control angry outputs Nanda et al. (2024). Additionally, other research demonstrates that SAEs can reconstruct vectors responsible for refusing to answer harmful questions as well as directions that produce sycophantic responses neverix et al. (2024). Please refer to Appendix E for more LLM applications.
6 Conclusions
In this survey, we provided a comprehensive examination of SAEs as a promising approach to interpreting and understanding LLMs. SAEs effectively address the challenge of polysemanticity through learning overcomplete, sparse representations that disentangle superimposed features into more interpretable units. We have systematically explored the foundational principles, technical frameworks, evaluation methodologies, and real-world applications of SAEs in the context of LLM analysis. While SAEs have demonstrated considerable success in revealing the internal mechanisms of LLMs, several challenges remain. In Appendix G, we discussed detailed research challenges, including the incompleteness of concept dictionaries, limited theoretical foundations, persistent reconstruction errors, and substantial computational requirements. Despite these challenges, SAEs continue to evolve through architectural innovations and improved training strategies, offering deeper insights into the inner workings of complex LLMs.
Limitations
Many concepts central to SAEs, such as polysemanticity, superposition, and feature disentanglement, have been extensively studied under different frameworks including distributed representations, disentangled representation learning, and sparse coding in computer vision and signal processing. This paper focuses specifically on introducing SAEs during the LLM era, examining their applications in interpreting transformer-based language models, and does not extensively cover these concepts’ history in adjacent fields.
Appendix A Why Sparse Autoencoders?
Explaining Individual Neurons.
The development of interpretability techniques for LLMs has progressed in stages rather than as a single step. From 2022 to 2023, researchers at OpenAI and Anthropic focused on understanding LLMs by examining individual neurons. OpenAI, for instance, leveraged GPT-4 to generate natural language explanations for neurons in models like GPT-2, attempting to map specific neuron activations to concrete linguistic or conceptual features Bills et al. (2023). Similarly, Anthropic built small toy models trained on synthetic data to observe how features are stored in neurons. These early investigations showed that analyzing single neurons can provide initial insights Elhage et al. (2022). In addition, it is worth noting that the study of explaining individual neurons by labeling interpretable features to them has been extensively explored in studies (Radford et al., 2017; Donnelly and Roegiest, 2019; Nguyen et al., 2019; Szegedy et al., 2013) prior to the introduction of mechanistic interpretability.
SAEs for Small-size Language Models.
SAEs for Large Language Models.
Appendix B Connection of SAEs to the Broader Field of Interpretability
The field of mechanistic interpretability (MI) has been critiqued for its insufficient engagement with the broader interpretability and NLP research literature Bereska and Gavves (2024a); Saphra and Wiegreffe (2024). Many of the research topics within MI, such as polysemanticity, superposition, and SAEs, have been investigated in prior and concurrent non-MI fields, often under different terminologies while addressing the same fundamental challenges (Saphra and Wiegreffe, 2024; Elhage et al., 2022). For instance, the study of polysemanticity and superposition, which aims to understand how features are encoded in the model activations, have been studied in the context of distributed representations (Hinton, 1984; Mikolov et al., 2013b, a; Arora et al., 2018; Olah, 2023), disentangled representations (Higgins et al., 2018; Kim and Mnih, 2018; Locatello et al., 2019), and concept-based interpretability (Nicolson et al., 2024; Kim et al., 2018). Similarly, SAEs are closely related to and draw inspiration from earlier lines of research on sparse coding and dictionary learning (Olshausen and Field, 1997; Gregor and LeCun, 2010; Faruqui et al., 2015; Subramanian et al., 2018). These methods, like SAEs, posit the feature sparsity hypothesis (Elhage et al., 2022) and aim to learn an overcomplete representation to disentangle the features from activation in superposition. Since these fields pursue similar goals or study the same research problems, the current disconnect causes issues such as missing relevant literature, hindering collaboration, unintentionally redefining established concepts, rediscovering existing techniques, and overlooking well-known baselines. Therefore, it is imperative for the MI community to bridge these gaps and more actively integrate findings from related non-MI research.
Appendix C Different SAE Variants
C.1 Improve Architecture